Let x be a continuous random variable that is normally distributed with a mean of 36 and a standard deviation of 5. Find the probability the x is less than 38. Round to four decimal places.
Answer #1

Let x be a continuous random variable that is normally distributed with a mean of 36 and a standard deviation of 5. Find the probability the x is less than 38.

Pr( x < 38 ) = Pr( z < 0.4) = 0.6554 Use Normal probability table and see 0.4 corresponding value in 0 colunm

Where z = (x - mean)/standard deviation = (38 - 36) / 5 = 0.4

OR

Pr( x < 38 ) = 0.6554 Use Excel command " =NORMDIST(38, 36, 5 ,TRUE) " , we get 0.6554